An 8.75 kg point mass and a 15.0 kg point mass are held in place 50.0 cm apart. A particle of mass m is released from a point between the two masses 18.0 cm from the 8.75 kg mass along the line connecting the two fixed masses.

1. Find the magnitude of the acceleration of the particle.

a= ________ m/s^2

Can you finish this problem before 12 pm central time cause around that time is when the assignment is due. Thank you for your time, have a great day!
Answer #1

Here,

let the acceleration of the particle is a

for the net force acting on the particle

6.673 *10^-11 * m*(15/0.32^2 - 8.75/0.18^2) = m * a

6.673 *10^-11 *(15/0.32^2 - 8.75/0.18^2) = a

solving for a

a = -8.24 *10^-9 m/s^2

the magnitude of acceleration is 8.24 *10-9 m/s^2